# Stabilizer Entropy and entanglement complexity in the Sachdev-Ye-Kitaev model

### v2

@@ Appendix B Capacity of entanglement - quick rederivation
For the convenience of the reader, we compute here Eq. 16 for the pure Haar random states. Original derivations were performed in [5, 15]. For Haar random pure states, the average Rényi entropy to leading order inN, with positive integer $\alpha$, is given by
@@
$\tilde{S}_{\alpha}^{\rm Haar}=-\alpha H_{\alpha-\frac{1}{2}}+\alpha H_{\alpha+1}+\frac{1}{2}(\alpha(N-4)\log(2)-\log(\pi))+\log\left(\frac{2^{\frac{1}{2}(-\alpha L+4\alpha+L)}\Gamma\left(\alpha+\frac{1}{2}\right)}{\Gamma(\alpha+2)}\right),$ (20)
@@ Appendix D Analytical derivation of the logarithmic anti-flatness for the SYK-2 model
$\overline{S}^{(3)}=-\dfrac{R}{4\pi f}\int\dfrac{\sqrt{p(1-p)+f(1-f)-\frac{1}{4}}}{p(1-p)}\ln{\left[p_{i}^{3}+(1-p_{i})^{3}\right]}$ (26)
@@
$\overline{S}^{(3)}=-\dfrac{R}{2\pi f}\int\limits_{\lambda_{-}}^{\lambda_{+}}\dfrac{\sqrt{\frac{1-\lambda^{2}}{4}+f(1-f)-\frac{1}{4}}}{1-\lambda^{2}}\ln{\left[\left(\dfrac{\lambda+1}{2}\right)^{3}+\left(1-\frac{\lambda+1}{2}\right)^{3}\right]}{\rm d}\lambda,$ (27)
@@
$\overline{S}^{(3)}=\dfrac{R}{4\pi f}\sum\limits_{n=1}^{\infty}\frac{1}{n}\left(\frac{3}{4}\right)^{n}\underbrace{\int\limits_{-2\sqrt{f(1-f)}}^{+2\sqrt{f(1-f)}}\dfrac{\sqrt{(1-\lambda^{2})+4f(1-f)-1}}{1-\lambda^{2}}\left(1-\lambda^{2}\right)^{n}{\rm d}\lambda}_{-2(f-1)f\pi\,_{2}F_{1}\left(\frac{1}{2},1-n,2,-4(f-1)f\right)}.$ (29)
Using the following definition of the logarithmic anti-flatness
@@
$\overline{F}(R,f)=2R(1-f)\sum\limits_{n=1}^{\infty}\dfrac{1}{n}\left(\dfrac{1}{2^{n}}-\dfrac{1}{2}\dfrac{3^{n}}{4^{n}}\right)\,_{2}F_{1}\left(\frac{1}{2},1-n\ ,2,4f(1-f)\right),$ (31)
In Fig. 8 we showcase the validity of the derived expression for the SYK-2 model.
@@ Appendix E Hamiltonian spectral statistics
Figure 9: Density of states (DOS) of the $H_{g}$ spectrum for $N=22$ fermions with support on $11$ qubits, with $M=40$ realizations of the disordered Hamiltonian. We considered $40960$ eigenvalues for each panel and corresponding $g$. See text for details and discussions. An interesting observation regarding the DOS support is highlighted below.
@@
We highlight a noteworthy aspect of the smooth change between the SYK4 and SYK2 DOS that can not be directly observed in Fig. 9. More specifically, for small values of $g$ in the range $g\in[0,0.1]$, we observe the shrinking of the energy support of the Hamiltonian eigenvalues and correspondingly the DOS support. We make this observation for all system sizes up to $N=22$ for which we computed the full spectrum. For values of $g$ that are larger than $0.1$, the support grows and the DOS starts spreading rather than shrinking. This change in behavior in the DOS could been related to our observations presented in Ref. 3 for the ground state. Interestingly, the ground state properties are sensitive to these changes in the Hamiltonian behavior, while the middle of the spectrum states are not. This phenomenology can potentially be related to the findings in [12] regarding the low-temperature behavior of the interpolated SYK model.
Figure 10: Spectral statistics results. Left panel: KL divergence (defined in the main text) between known Gaussian ensemble RMT distributions and the underlying dataset for the consecutive gap ratios PDF $P(r)$. Right panel: The averaged consecutive gap ratios. We considered $40960$ eigenvalues for each panel and corresponding $g$. Smaller systems size exhibit similar findings to the presented results for $N=22$ in this figure.
@@
$\bar{r}=\left\langle\left\langle\frac{\min(s_{k,j},s_{k+1,j})}{\max(s_{k,j},s_{k+1,j})}\right\rangle\right\rangle_{2^{N/2-1}-2,M}=\dfrac{1}{M(2^{N/2-1}-2)}\sum\limits_{j=1}^{M}\sum\limits_{k=1}^{2^{N/2-1}-2}\frac{\min(s_{k,j},s_{k+1,j})}{\max(s_{k,j},s_{k+1,j})}$ (35)
@@ Appendix F Analysis of the gap
In Fig. 11, we show the behavior of the average ground state to the first excited state gap for different choices of the interpolation parameter $g$. Fitting a power law function to the data we observe consistency with the expected $N^{-1}$ power law with exponent $-1$ as $g$ approaches 1. On the other hand, when we approach the SYK4 point, i.e. $g$ going to 0, we are still able to fit a power law function with reasonable precision. This implies that for the finite system data, it is hard to differentiate between a power-law and the exponential behavior expected for the SYK4 model, indicating the ineffectiveness of this particular probe.
@@ Appendix G Entanglement Spectral Statistics (ESS) and normalized Reduced Density Matrix (RDM) spectrum
In this section, we complete our survey of the ESS of the middle spectrum state not presented in the main text. More specifically, we repeat the same analysis presented in Fig. 4, but for the middle of the spectrum state (MS) at $E\leavevmode\nobreak\ \sim 0$, see Fig. 9. In Fig. 12, we observe that for the considered system sizes the ESS does not show agreement with RMT universal ensembles and shows good agreement with the uncorrelated Poisson distribution.
In Fig. 13, we expand on this point and demonstrate a broader view-point of the absence of universal behavior in the RDM eigenvalues. In the top panel, we show the results for pure Haar random states on $N=11$ qubits. To generate such states we simply draw real and imaginary parts of the expansion coefficients of $|\psi\rangle$ as random Gaussian variables, and normalize the output vector state [20]. This procedure provides an excellent agreement with the predicted Marchenko-Pasture (M-P) distribution defined in the main text and represented with the purple line in the middle panels. We note that the disagreement between obtained values and the M-P distribution comes from the odd/even effect that vanishes in the thermodynamic limit. The final panel of the upper row shows that ESS for the Haar states does conform with the RMT prediction and universal GUE ensemble. The Haar example serves as a reference point to the results we obtain for the MS state of the interpolated model. More specifically, we can observe that the RDM spectrum itself shows quite similar features to that of a Haar random state, however, the ESS point to an uncorrelated ensemble and the non-universal Poisson distribution. These unorthodox features of SYK4 and of the interpolated model, when it comes to universality statistics probes, are most likely part of the broader context of the particular structure that the eigenstates entail and are evidenced already in the above mentioned chapter on the Hamiltonian spectrum statistics.
Figure 13: Haar vs middle of the spectrum (MS) SYK states comparative analysis across a single bipartition (same results for other bipartitions). Upper row: Haar states included by influenced by the odd/een effects (the reason for the mismatch with the purple M-P distribution [19]); Middle row: MS state for the $g=0$ SYK4 Hamiltonian; Lower row: MS state for $g=0.01$. Other parameters: number of realization $M=100$, nb of spins $N=11$ (fermions $N=22$) and across bipartition $[1,2,3,4,5]$.

### v3

@@ Appendix B Capacity of entanglement - quick rederivation
For the convenience of the reader, we compute here Eq. (17) for the pure Haar random states. Original derivations were performed in [86, 87]. For Haar random pure states, the average Rényi entropy to leading order in $N$, with positive integer $\alpha$, is given by
@@
$\tilde{S}_{\alpha}^{\rm Haar}$ $=-\alpha H_{\alpha-\frac{1}{2}}+\alpha H_{\alpha+1}+\frac{1}{2}(\alpha(N-4)\log(2)-\log(\pi))$
$+\log\left(\frac{2^{\frac{1}{2}(-\alpha L+4\alpha+L)}\Gamma\left(\alpha+\frac{1}{2}\right)}{\Gamma(\alpha+2)}\right),$ (26)
@@ Appendix D Analytical derivation of the logarithmic anti-flatness for the SYK-2 model
Another measure of anti-flatness that is a numerically and analytically accessible quantity is the logarithmic anti-flatness [64]
$\mathrm{F}(\rho_{R}):=2\left(S_{2}(\rho_{R})-S_{3}(\rho_{R})\right).$ (29)
@@
$\overline{S}^{(3)}$ $=-\dfrac{R}{4\pi f}\int\dfrac{\sqrt{p(1-p)+f(1-f)-\frac{1}{4}}}{p(1-p)}$
×$\times\ln{\left[p_{i}^{3}+(1-p_{i})^{3}\right]}$ (33)
@@
$\overline{S}^{(3)}=-\dfrac{R}{2\pi f}\int\limits_{\lambda_{-}}^{\lambda_{+}}\dfrac{\sqrt{\frac{1-\lambda^{2}}{4}+f(1-f)-\frac{1}{4}}}{1-\lambda^{2}}$
×$\times\ln{\left[\left(\dfrac{\lambda+1}{2}\right)^{3}+\left(1-\frac{\lambda+1}{2}\right)^{3}\right]}{\rm d}\lambda,$ (34)
@@
$\overline{S}^{(3)}$ $=\dfrac{R}{4\pi f}\sum\limits_{n=1}^{\infty}\frac{1}{n}\left(\frac{3}{4}\right)^{n}\int\limits_{-2\sqrt{f(1-f)}}^{+2\sqrt{f(1-f)}}\left(1-\lambda^{2}\right)^{n}$
×$\times\dfrac{\sqrt{(1-\lambda^{2})+4f(1-f)-1}}{1-\lambda^{2}}{\rm d}\lambda.$ (36)
The integrable can be evaluate exactly to yield $-2(f-1)f\pi\,_{2}F_{1}\left(\frac{1}{2},1-n,2,-4(f-1)f\right)$. Using the following definition of the logarithmic anti-flatness
@@
$\overline{F}(R,f)=$ $2R(1-f)\sum\limits_{n=1}^{\infty}\dfrac{1}{n}\left(\dfrac{1}{2^{n}}-\dfrac{1}{2}\dfrac{3^{n}}{4^{n}}\right)\,$
×$\times_{2}F_{1}\left(\frac{1}{2},1-n\ ,2,4f(1-f)\right),$ (38)
In Fig. 9 we showcase the validity of the derived expression for the SYK-2 model. For Haar random state in the large $N$ limit the logarithmic anti-flatness approach system-size-independent value of $\mathrm{F}^{\mathrm{Haar}}=\log{(5/4)}\approx 0.223$ [64].
@@ Appendix E Hamiltonian spectral statistics
Figure 10: Density of states (DOS) of the $H_{g}$ spectrum for $N=22$ fermions with support on 111111 qubits, with $M=40$ realizations of the disordered Hamiltonian. We considered 409604096040960 eigenvalues for each panel and corresponding $g$. See text for details and discussions. An interesting observation regarding the DOS support is highlighted below.
@@
We highlight a noteworthy aspect of the smooth change between the SYK4 and SYK2 DOS that can not be directly observed in Fig. 10. More specifically, for small values of $g$ in the range $g\in[0,0.1]$, we observe the shrinking of the energy support of the Hamiltonian eigenvalues and correspondingly the DOS support. We make this observation for all system sizes up to $N=22$ for which we computed the full spectrum. For values of $g$ that are larger than 0.10.10.1, the support grows and the DOS starts spreading rather than shrinking. This change in behavior in the DOS could been related to our observations presented in Ref. (3) for the ground state. Interestingly, the ground state properties are sensitive to these changes in the Hamiltonian behavior, while the middle of the spectrum states are not. This phenomenology can potentially be related to the findings in [14] regarding the low-temperature behavior of the interpolated SYK model.
@@
Figure 11: Spectral statistics results. Left panel: KL divergence (defined in the main text) between known Gaussian ensemble RMT distributions and the underlying dataset for the consecutive gap ratios PDF $P(r)$. Right panel: The averaged consecutive gap ratios. We considered 409604096040960 eigenvalues for each panel and corresponding $g$. Smaller systems size exhibit similar findings to the presented results for $N=22$ in this figure.
@@
$\bar{r}$ $=\left\langle\left\langle\frac{\min(s_{k,j},s_{k+1,j})}{\max(s_{k,j},s_{k+1,j})}\right\rangle\right\rangle_{2^{N/2-1}-2,M}$
$=\dfrac{1}{M(2^{N/2-1}-2)}\sum\limits_{j=1}^{M}\sum\limits_{k=1}^{2^{N/2-1}-2}\frac{\min(s_{k,j},s_{k+1,j})}{\max(s_{k,j},s_{k+1,j})}$ (42)
@@ Appendix F Analysis of the gap
In Fig. 12, we show the behavior of the average ground state to the first excited state gap for different choices of the interpolation parameter $g$. Fitting a power law function to the data we observe consistency with the expected $N^{-1}$ power law with exponent 1-1- 1 as $g$ approaches 1. On the other hand, when we approach the SYK4 point, i.e. $g$ going to 0, we are still able to fit a power law function with reasonable precision. This implies that for the finite system data, it is hard to differentiate between a power-law and the exponential behavior expected for the SYK4 model, indicating the ineffectiveness of this particular probe.
Figure 14: Haar vs middle of the spectrum (MS) SYK states comparative analysis across a single bipartition (same results for other bipartitions). Upper row: Haar states included by influenced by the odd/een effects (the reason for the mismatch with the purple M-P distribution [41]); Middle row: MS state for the $g=0$ SYK4 Hamiltonian; Lower row: MS state for $g=0.01$. Other parameters: number of realization $M=100$, nb of spins $N=11$ (fermions $N=22$) and across bipartition [1,2,3,4,5][1,2,3,4,5][ 1 , 2 , 3 , 4 , 5 ].
@@ Appendix G Entanglement Spectral Statistics (ESS) and normalized Reduced Density Matrix (RDM) spectrum
In this section, we complete our survey of the ESS of the middle spectrum state not presented in the main text. More specifically, we repeat the same analysis presented in Fig. (4), but for the middle of the spectrum state (MS) at $E\leavevmode\nobreak\ \sim 0$, see Fig. 10. In Fig. 13, we observe that for the considered system sizes the ESS does not show agreement with RMT universal ensembles and shows good agreement with the uncorrelated Poisson distribution.
In Fig. 14, we expand on this point and demonstrate a broader view-point of the absence of agreement between Haar expectation in the RDM eigenvalues. In the top panel, we show the results for pure Haar random states on $N=11$ qubits. To generate such states we simply draw real and imaginary parts of the expansion coefficients of $|\psi\rangle$ as random Gaussian variables, and normalize the output vector state [57]. This procedure provides an excellent agreement with the predicted Marchenko-Pasture (M-P) distribution defined in the main text and represented with the purple line in the middle panels. We note that the disagreement between obtained values and the M-P distribution comes from the odd/even effect that vanishes in the thermodynamic limit. The final panel of the upper row shows that ESS for the Haar states does conform with the RMT prediction and universal GUE ensemble. The Haar example serves as a reference point to the results we obtain for the MS state of the interpolated model. More specifically, we can observe that the RDM spectrum itself shows quite similar features to that of a Haar random state, however, the ESS point to an uncorrelated ensemble and the non-universal Poisson distribution. These unorthodox features of SYK4 and of the interpolated model, when it comes to universality statistics probes, are most likely part of the broader context of the particular structure that the eigenstates entail and are evidenced already in the above mentioned section on the Hamiltonian spectrum statistics.
Figure 15: Ensemble-averaged SRE $\langle\mathcal{M}_{2}\rangle_{M}$ in middle-of-the-spectrum (MS) states of the interpolated Hamiltonian $H(g)$. Inset: total SRE for a fixed interpolation parameter $g=0.3$ as a function of system size $N$.
Appendix H Stabilizer Rényi Entropy of Middle of the spectrum states
We present results for the Stabilizer Rényi entropy (SRE) measured in middle-of-the-spectrum eigenstates of the interpolated SYK Hamiltonian $H(g)$. As shown in Fig. 15, the behavior of the SRE qualitatively resembles that of the entanglement entropy for these states: the SYK-4 phase dominates across most values of the interpolation parameter $g$. In the inset of the figure, we show how the SRE changes with system size for value $g=0.3$, which is deep inside the SYK-4 regime and does not allow for symmetry constraints present for the pure $g=0$ SYK-4 model. In particular, in terms of non-stabilizerness, also high-energy states fail to reach Haar-type universality.